### v2

@@ 2.1 Nonconvex-strongly Concave Setting
We first propose a completely parameter-free AGP algorithm for solving nonconvex-strongly concave minimax problems. More specifically, in order to estimate the Lipschitz constants for $\nabla_{x}f(\cdot,y)$, $\nabla_{y}f(\cdot,y)$, $\nabla_{y}f(x,\cdot)$ denoted as $L_{11}$, $L_{12}$ and $L_{22}$, respectively, and the strongly concave constant $\mu$, at each iteration of the proposed algorithm, we aim to find a tuple $(l_{11}^{k,i},l_{12}^{k,i},l_{22}^{k,i},\mu_{k,i})$ by backtracking such that the following conditions for $(x_{k,i},y_{k,i})$ are satisfied:
$C_{1}^{k,i}$ $=f(x_{k,i},y_{k})-f(x_{k},y_{k})-\langle\nabla_{x}f(x_{k},y_{k}),x_{k,i}-x_{k}\rangle-\frac{l^{k,i}_{11}}{2}\|x_{k,i}-x_{k}\|^{2}\leq 0,$ (C1)
@@
$C_{3}^{k,i}$ $=\|\nabla_{y}f(x_{k,i},y_{k,i})-\nabla_{y}f(x_{k,i},y_{k})\|-l_{22}^{k,i}\|y_{k,i}-y_{k}\|\leq 0,$ (C3)
@@
  Step 1: Input $x_{1},y_{1}$, $\beta_{1}$, $\gamma_{1}$, $l_{11}^{1}$, $l_{12}^{1}$, $l_{22}^{1}$, $\mu_{1}$; Set $k=1$.
  Step 2: Updata $x_{k}$ and $y_{k}$:
   (a): Set $i=1$, $l_{11}^{k,i}=l_{11}^{k}$, $l_{12}^{k,i}=l_{12}^{k}$, $l_{22}^{k,i}=l_{22}^{k}$, $\mu_{k,i}=\mu_{k}$, $\beta_{k,i}=\beta_{k}$, $\gamma_{k,i}=\gamma_{k}$.
@@
$\beta_{k,i}={l_{11}^{k,i}}+\frac{l_{12}^{k,i}\mu_{k-1,i}}{8(l_{22}^{k-1,i})^{2}}+\frac{32(l_{12}^{k,i})^{2}(l_{22}^{k-1,i})^{2}}{\mu_{k,i}\mu_{k-1,i}^{2}},\quad\gamma_{k,i}=\frac{8(l_{22}^{k,i})^{2}}{\mu_{k,i}},$
@@
It is worth noting that the parameter-free NeAda (Yang et al., 2022) and TiAda (Li et al., 2022) algorithms with adaptive step size are only used to solve the minimax problems when $x$ is unconstrained, while the proposed PF-AGP-NSC algorithm can solve the constrained case. Compared with the SGDA-B (Xu et al., 2023) algorithm and AGDA+ (Zhang et al., 2024) algorithm using the backtracking framework under the nonconvex-strongly concave setting, AGDA+ and SGDA-B still require knowledge of the strongly concave constant $\mu$, while the proposed PF-AGP-NSC algorithm is a completely parameter-free algorithm that does not require any prior knowledge of the parameters.
@@ 2.1.1 Complexity analysis
$\nabla\tilde{G}(x_{k},y_{k})=\begin{bmatrix}\beta_{k}\left(x_{k}-P_{x}\left(x_{k}-\frac{1}{\beta_{k}}\nabla_{x}f_{k}(x_{k},y_{k})\right)\right)\\ \gamma_{k}\left(y_{k}-P_{y}\left(y_{k}+\frac{1}{\gamma_{k}}\nabla_{y}f_{k}(x_{k},y_{k})\right)\right)\end{bmatrix}.$
@@
The backtracking strategy (C1) implies that
@@
Next, we provide an estimate of upper bound of the difference between $f(x_{k+1},y_{k+1})$ and $f(x_{k},y_{k})$. First, we need to make the following assumption on the parameter $\gamma_{k}$.
@@
Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 1, then we have
@@
$\quad\leq-\left(\beta_{k}-\frac{l_{11}^{k}}{2}-\frac{l_{12}^{k}}{2\gamma_{k-1}}\right)\|x_{k+1}-x_{k}\|^{2}+{\gamma_{k-1}}\|y_{k+1}-y_{k}\|^{2}$
$\quad\quad-\left(\mu_{k-1}-\frac{\gamma_{k-1}}{2}-\frac{(l_{22}^{k-1})^{2}}{2\gamma_{k-1}}\right)\|y_{k}-y_{k-1}\|^{2}.$ (11)
@@
$=$ $\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle+\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),v_{k+1}\rangle$
@@
Next, we estimate the three terms on the right-hand side of (14) respectively. By the backtracking strategy (C2), (C3) and the Cauchy-Schwarz inequality, we have
$\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle\leq\frac{l_{12}^{k}}{2\gamma_{k}}\|x_{k+1}-x_{k}\|^{2}+\frac{\gamma_{k}}{2}\|y_{k+1}-y_{k}\|^{2}$ (15)
@@
$\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),v_{k+1}\rangle\leq\frac{(l_{22}^{k-1})^{2}}{2\gamma_{k-1}}\|y_{k}-y_{k-1}\|^{2}+\frac{\gamma_{k-1}}{2}\|v_{k+1}\|^{2}.$ (16)
By the backtracking strategy (C4) of $\mu_{k-1}$, we obtain
$\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle\leq-\mu_{k-1}\|y_{k}-y_{k-1}\|^{2}.$ (17)
@@
$f(x_{k+1},y_{k+1})-f(x_{k+1},y_{k})$ $\leq\frac{l_{12}^{k}}{2\gamma_{k}}\|x_{k+1}-x_{k}\|^{2}-\left(\mu_{k-1}-\frac{\gamma_{k-1}}{2}-\frac{(l_{22}^{k-1})^{2}}{2\gamma_{k-1}}\right)\|y_{k}-y_{k-1}\|^{2}$
$\quad+\left(\frac{\gamma_{k}}{2}+\frac{\gamma_{k-1}}{2}\right)\|y_{k+1}-y_{k}\|^{2}.$ (19)
@@
Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 1, and denoting
$f_{k+1}$ $=f(x_{k+1},y_{k+1}),\quad S_{k+1}=\frac{2\gamma^{2}_{k}}{\mu_{k}}\|y_{k+1}-y_{k}\|^{2},$
$F_{k+1}$ $=f_{k+1}+S_{k+1}-\left(\mu_{k}+\frac{7\gamma_{k}}{2}-\frac{(l^{k}_{22})^{2}}{2\gamma_{k}}-\frac{2(l^{k}_{22})^{2}}{\mu_{k}}\right)\|y_{k+1}-y_{k}\|^{2}-\frac{2\gamma_{k}^{2}}{\mu_{k}}\mathcal{D}^{2}_{y},$ (21)
then we have
$F_{k+1}-F_{k}\leq$ $-\left(\frac{2\mu_{k}\gamma_{k}-(l_{22}^{k})^{2}}{2\gamma_{k}}+\frac{\mu_{k}\gamma_{k}-4(l_{22}^{k})^{2}}{2\mu_{k}}\right)\|y_{k+1}-y_{k}\|^{2}$
$\quad-\left(\beta_{k}-\frac{l_{11}^{k}}{2}-\frac{l_{12}^{k}}{2\gamma_{k-1}}-\frac{2(l_{12}^{k})^{2}\gamma_{k-1}}{\mu_{k-1}\mu_{k}}\right)\|x_{k+1}-x_{k}\|^{2},$ (22)
@@
$\langle\nabla_{y}f(x_{k},y_{k-1})-\nabla_{y}f(x_{k+1},y_{k})+\gamma_{k}(y_{k+1}-y_{k})-\gamma_{k-1}(y_{k}-y_{k-1}),y_{k+1}-y_{k}\rangle\leq 0,$ (24)
@@
$\gamma_{k}\|y_{k+1}-y_{k}\|^{2}-\gamma_{k-1}\langle y_{k}-y_{k-1},y_{k+1}-y_{k}\rangle\leq\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle.$ (25)
@@
$=$ $\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle+\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),v_{k+1}\rangle$
@@
$\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle\leq\frac{(l_{12}^{k})^{2}}{2\mu_{k}}\|x_{k+1}-x_{k}\|^{2}+\frac{\mu_{k}}{2}\|y_{k+1}-y_{k}\|^{2}.$ (28)
@@
$\frac{\gamma_{k}}{2}\|y_{k+1}-y_{k}\|^{2}-\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}$ $\leq\frac{(l_{12}^{k})^{2}}{2\mu_{k}}\|x_{k+1}-x_{k}\|^{2}+\left(\frac{\mu_{k}}{2}+\frac{\gamma_{k-1}}{2}-\frac{\gamma_{k}}{2}\right)\|y_{k+1}-y_{k}\|^{2}$
$\quad-\left(\mu_{k-1}-\frac{(l_{22}^{k-1})^{2}}{2\gamma_{k-1}}\right)\|y_{k}-y_{k-1}\|^{2}.$ (29)
By multiplying $\frac{4\gamma_{k-1}}{\mu_{k-1}}$ on both sides of (29), using the definition of $S_{k+1}$ and Assumption 2, we obtain
$\begin{split}S_{k+1}-S_{k}&\leq\frac{2(l_{12}^{k})^{2}\gamma_{k-1}}{\mu_{k-1}\mu_{k}}\|x_{k+1}-x_{k}\|^{2}+2\gamma_{k-1}\|y_{k+1}-y_{k}\|^{2}\\ &\quad-\left({4\gamma_{k-1}}-\frac{2(l_{22}^{k-1})^{2}}{\mu_{k-1}}\right)\|y_{k}-y_{k-1}\|^{2}+\left(\frac{2\gamma_{k}^{2}}{\mu_{k}}-\frac{2\gamma^{2}_{k-1}}{\mu_{k-1}}\right)\mathcal{D}_{y}^{2}.\end{split}$
@@
Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 1, $\beta_{k}={l_{11}^{k}}+\frac{l_{12}^{k}\mu_{k-1}}{8(l_{22}^{k-1})^{2}}+\frac{32(l_{12}^{k})^{2}(l_{22}^{k-1})^{2}}{\mu_{k}\mu_{k-1}^{2}}$, $\gamma_{k}=\frac{8(l_{22}^{k})^{2}}{\mu_{k}}$, then it holds that
@@
where $F_{2}$ is defined in (21), $\underline{F}=\underline{f}-\left(\frac{224L_{22}^{2}}{\mu}+\mu_{1}-\frac{2(l_{22}^{1})^{2}}{\mu_{1}}-\frac{\left(l_{22}^{1}\mu\right)^{2}}{128L_{22}^{2}}+\frac{2^{13}L_{22}^{4}}{\mu^{3}}\right)\mathcal{D}^{2}_{y}$, with $\underline{f}=\min_{(x,y)\in X\times Y}f(x,y)$, $\mathcal{D}_{y}=\max\{\|y_{1}-y_{2}\|\mid y_{1},y_{2}\in\mathcal{Y}\}$, $d_{1}=\min\{\frac{\beta_{\text{min}}}{2\beta_{\text{max}}^{2}+4L^{2}_{12}},\frac{15\mu^{3}+2^{7}l_{22}^{1}\mu}{2^{15}L_{22}^{4}}\}$, $\beta_{\text{min}}={l^{1}_{11}}+\frac{l^{1}_{12}\mu}{64(L_{22})^{2}}+\frac{32(l_{12}^{1})^{2}(l^{1}_{22})^{2}}{\mu_{1}^{3}}$ and $\beta_{\text{max}}=2L_{11}+\frac{\mu_{1}L_{12}}{4(l^{1}_{22})^{2}}+\frac{4096(L_{12})^{2}(L_{22})^{2}}{\mu^{3}}$.
@@
Let $d^{(k)}_{1}={\min\left\{\frac{\frac{2\mu_{k}\gamma_{k}-\left(l^{k}_{22}\right)^{2}}{2\gamma_{k}}+\frac{\mu_{k}\gamma_{k}-4\left(l_{22}^{k}\right)^{2}}{2\mu_{k}}}{2\gamma_{k}^{2}},\frac{\beta_{k}-\frac{l^{k}_{11}}{2}-\frac{l^{k}_{12}}{2\gamma_{k-1}}-\frac{2(l^{k}_{12})^{2}\gamma_{k-1}}{\mu_{k-1}\mu_{k}}}{\beta_{k}^{2}+2L_{12}^{2}}\right\}}>0$. Multiplying both sides of (34) by $d^{(k)}_{1}$, and using (22) in Lemma 5, we have
@@
$F_{T(\varepsilon)+1}$ $=f_{T(\varepsilon)+1}+S_{T(\varepsilon)+1}-\frac{2\gamma_{T(\varepsilon)}^{2}}{\mu_{T(\varepsilon)}}\mathcal{D}^{2}_{y}$
$\quad-\left(\frac{7}{2}\gamma_{T(\varepsilon)+1}+\mu_{T(\varepsilon)+1}-\frac{2(l_{22}^{T(\varepsilon)+1})^{2}}{\mu_{T(\varepsilon)+1}}-\frac{(l_{22}^{T(\varepsilon)+1})^{2}}{2\gamma_{T(\varepsilon)+1}}\right)||y_{T(\varepsilon)+1}-y_{T(\varepsilon)}||^{2}$
$\geq\underline{f}-\left(\frac{224L_{22}^{2}}{\mu}+\mu_{1}-\frac{2(l_{22}^{1})^{2}}{\mu_{1}}-\frac{\left(l_{22}^{1}\mu\right)^{2}}{128L_{22}^{2}}+\frac{2^{13}L_{22}^{4}}{\mu^{3}}\right)\mathcal{D}^{2}_{y}$
$=\underline{F},$ (38)
@@
At the $k$-th iteration of Algorithm 1, the number of backtracking step is upper bounded by $\log_{2}\left(\frac{2L_{11}}{l_{11}^{1}}\right)+\log_{2}\left(\frac{2L_{12}}{l_{12}^{1}}\right)+\log_{2}\left(\frac{2L_{22}}{l_{22}^{1}}\right)+\log_{2}\left(\frac{2\mu_{1}}{\mu}\right)=\mathcal{O}(1)$. By Theorem 6, the total number of gradient calls of Algorithm 1 to obtain an $\varepsilon$-stationary point that satisfies $\|\nabla G(x_{k},y_{k})\|\leq\varepsilon$ is upper bounded by $\mathcal{O}(L\kappa^{3}\varepsilon^{-2})$ under the nonconvex-strongly concave setting, where $\kappa$ is the condition number.

### v3

@@ 2.1 Nonconvex-strongly Concave Setting
We first propose a completely parameter-free AGP algorithm for solving nonconvex-strongly concave minimax problems. More specifically, in order to estimate the Lipschitz constants for $\nabla_{x}f(\cdot,y)$, $\nabla_{y}f(\cdot,y)$, $\nabla_{y}f(x,\cdot)$ denoted as $L_{11}$, $L_{12}$ and $L_{22}$, respectively, and the strongly concavity modulus $\mu$, at each iteration of the proposed algorithm, we aim to find a tuple $(l_{11}^{k,i},l_{12}^{k,i},l_{22}^{k,i},\mu_{k,i})$ by backtracking such that the following conditions for $(x_{k,i},y_{k,i})$ are satisfied:
$C_{1}^{k,i}=$ $f(x_{k,i},y_{k})-f(x_{k},y_{k})-\langle\nabla_{x}f(x_{k},y_{k}),x_{k,i}-x_{k}\rangle-\frac{l^{k,i}_{11}}{2}\|x_{k,i}-x_{k}\|^{2}\leq 0,$
@@
$C_{3}^{k,i}=$ $l_{22}^{k,i}\langle\nabla_{y}f(x_{k,i},y_{k,i})-\nabla_{y}f(x_{k,i},y_{k,i}),y_{k,i}-y_{k}\rangle$
$+\|\nabla_{y}f(x_{k,i},y_{k,i})-\nabla_{y}f(x_{k,i},y_{k})\|^{2}\leq 0,$ (C3)
@@
  Step 1: Input $x_{1},y_{1}$, $\beta_{0}$, $\gamma_{0}$, $l_{11}^{0}$, $l_{12}^{0}$, $l_{22}^{0}$, $\mu_{0}$; Set $k=1$.
  Step 2: Update $x_{k}$ and $y_{k}$:
   (a): Set $i=0$, $l_{11}^{k,i}={l_{11}^{k-1}}{}$, $l_{12}^{k,i}={l_{12}^{k-1}}$, $l_{22}^{k,i}=l_{22}^{k-1}$, $\mu_{k,i}=\mu_{k-1}$, $\beta_{k,i}=\beta_{k-1}$, $\gamma_{k,i}=\gamma_{k-1}$.
@@
$\beta_{k,i}={l_{11}^{k,i}}+l_{12}^{k,i}+\frac{32(l_{12}^{k,i})^{2}(l^{k-1}_{12}+l^{k-1}_{22})}{\mu_{k,i}\mu_{k-1}},\quad\gamma_{k,i}=l_{12}^{k,i}+l_{22}^{k,i},$
@@
It is worth noting that the parameter-free NeAda (Yang et al., 2022) and TiAda (Li et al., 2022) algorithms with adaptive step size are only used to solve the minimax problems when $x$ is unconstrained, while the proposed PF-AGP-NSC algorithm can solve the constrained case. Compared with the SGDA-B (Xu et al., 2024) algorithm and AGDA+ (Zhang et al., 2025) algorithm using the backtracking framework under the nonconvex-strongly concave setting, SGDA-B still requires knowledge of the strongly concavity modulus $\mu$, while the proposed PF-AGP-NSC algorithm is a completely parameter-free algorithm that does not require any prior knowledge of the parameters. Moreover, the complexity of the proposed PF-AGP-NSC is better than AGDA+.
@@ 2.1.1 Complexity analysis
$\nabla\tilde{G}(x_{k},y_{k})=\begin{bmatrix}\beta_{k}\left(x_{k}-P_{\mathcal{X}}\left(x_{k}-\frac{1}{\beta_{k}}\nabla_{x}f_{k}(x_{k},y_{k})\right)\right)\\ \gamma_{k}\left(y_{k}-P_{\mathcal{Y}}\left(y_{k}+\frac{1}{\gamma_{k}}\nabla_{y}f_{k}(x_{k},y_{k})\right)\right)\end{bmatrix}.$
@@
The backtracking strategy (2.1) implies that
@@
Next, we provide an upper bound of the difference between $f(x_{k+1},y_{k+1})$ and $f(x_{k},y_{k})$. First, we need to make the following assumption on the parameter $\gamma_{k}$.
@@
Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 1,  if $\gamma_{k}\geq l_{22}^{k}$, then we have
@@
$\leq$ $-\left(\beta_{k}-\frac{l_{11}^{k}}{2}-\frac{l_{12}^{k}}{2}\right)\|x_{k+1}-x_{k}\|^{2}+\left(\frac{l^{k}_{12}}{2}+\frac{\gamma_{k-1}}{2}\right)\|y_{k+1}-y_{k}\|^{2}$
$-\left(\frac{\mu_{k-1}}{2}-\frac{\gamma_{k-1}}{2}\right)\|y_{k}-y_{k-1}\|^{2}.$ (11)
@@
$=$ $\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle$
$+\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),v_{k+1}\rangle$
@@
Next, we estimate the three terms on the right-hand side of (14) respectively. By the backtracking strategy (C2) and the Cauchy-Schwarz inequality, we have
$\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle\leq\frac{l_{12}^{k}}{2}\|x_{k+1}-x_{k}\|^{2}+\frac{l^{k}_{12}}{2}\|y_{k+1}-y_{k}\|^{2}$ (15)
@@
$\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),v_{k+1}\rangle\leq\frac{1}{2\gamma_{k-1}}\|f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1})\|+\frac{\gamma_{k-1}}{2}\|v_{k+1}\|.$ (16)
By the backtracking strategy (C3) and (C4), we obtain
$\begin{split}&\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle\\ \leq&-\frac{1}{2l_{22}^{k-1}}\|f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1})\|-\frac{\mu_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}.\end{split}$ (17)
@@
$f(x_{k+1},y_{k+1})-f(x_{k+1},y_{k})$
$\leq$ $\frac{l^{k}_{12}}{2}\|x_{k+1}-x_{k}\|^{2}+\left(\frac{l^{k}_{12}}{2}+\frac{\gamma_{k-1}}{2}\right)\|y_{k}-y_{k-1}\|^{2}-\left(\frac{\mu_{k-1}}{2}-\frac{\gamma_{k-1}}{2}\right)\|y_{k+1}-y_{k}\|^{2}.$ (19)
@@
Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 1, and denote
$f_{k+1}$ $=f(x_{k+1},y_{k+1}),\quad S_{k+1}=\frac{4\gamma^{2}_{k}}{\mu_{k}}\|y_{k+1}-y_{k}\|^{2},$
$F_{k+1}$ $=f_{k+1}+S_{k+1}-\frac{7\gamma_{k}}{2}\|y_{k+1}-y_{k}\|^{2}-\frac{4\gamma_{k}^{2}}{\mu_{k}}\mathcal{D}^{2}_{y}.$ (21)
If $\gamma_{k}\geq l_{22}^{k}$, then we have
$F_{k+1}-F_{k}\leq$ $-\left(2\gamma_{k}-\frac{l_{12}^{k}}{2}\right)\|y_{k+1}-y_{k}\|^{2}$
$-\left(\beta_{k}-\frac{l_{11}^{k}}{2}-\frac{l_{12}^{k}}{2}-\frac{16(l_{12}^{k})^{2}\gamma_{k-1}}{\mu_{k-1}\mu_{k}}\right)\|x_{k+1}-x_{k}\|^{2},$ (22)
@@
$\langle\nabla_{y}f(x_{k},y_{k-1})-\nabla_{y}f(x_{k+1},y_{k})+\gamma_{k}(y_{k+1}-y_{k})-\gamma_{k-1}(y_{k}-y_{k-1}),y_{k+1}-y_{k}\rangle$
$\leq 0,$ (24)
@@
$\gamma_{k}\|y_{k+1}-y_{k}\|^{2}-\gamma_{k-1}\langle y_{k}-y_{k-1},y_{k+1}-y_{k}\rangle$
$\leq\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle.$ (25)
@@
$=$ $\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle$
$+\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),v_{k+1}\rangle$
@@
$\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle$
$\leq$ $\frac{2(l_{12}^{k})^{2}}{\mu_{k}}\|x_{k+1}-x_{k}\|^{2}+\frac{\mu_{k}}{8}\|y_{k+1}-y_{k}\|^{2}.$ (28)
@@
$\frac{\gamma_{k}}{2}\|y_{k+1}-y_{k}\|^{2}-\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}$
$\leq$ $\frac{2(l_{12}^{k})^{2}}{\mu_{k}}\|x_{k+1}-x_{k}\|^{2}+\left(\frac{\mu_{k}}{8}+\frac{\gamma_{k-1}}{2}-\frac{\gamma_{k}}{2}\right)\|y_{k+1}-y_{k}\|^{2}$
$-\frac{\mu_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}.$ (29)
By multiplying  $\frac{8\gamma_{k-1}}{\mu_{k-1}}$ on both sides of (29), using the definition of $S_{k+1}$ and Assumption 2, we obtain
$\begin{split}S_{k+1}-S_{k}&\leq\frac{16(l_{12}^{k})^{2}\gamma_{k-1}}{\mu_{k-1}\mu_{k}}\|x_{k+1}-x_{k}\|^{2}+\gamma_{k-1}\|y_{k+1}-y_{k}\|^{2}\\ &\quad-{4\gamma_{k-1}}\|y_{k}-y_{k-1}\|^{2}+\left(\frac{4\gamma_{k}^{2}}{\mu_{k}}-\frac{4\gamma^{2}_{k-1}}{\mu_{k-1}}\right)\mathcal{D}_{y}^{2}.\end{split}$
@@
Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 1, $\beta_{k}={l_{11}^{k}}+l_{12}^{k}+\frac{32(l_{12}^{k})^{2}(l^{k-1}_{12}+l_{22}^{k-1})}{\mu_{k}\mu_{k-1}},\gamma_{k}=l_{12}^{k}+l_{22}^{k}$, then it holds that
@@
where $F_{2}$ is defined in (21), $\underline{F}=\underline{f}-\left(14L_{12}+\frac{128L_{12}}{\mu}\right)\mathcal{D}^{2}_{y}$, with $\mathcal{D}_{y}=\max\{\|y_{1}-y_{2}\|\mid y_{1},y_{2}\in\mathcal{Y}\}$, $d_{1}=\min\{\frac{3}{16l^{1}_{12}},\frac{\beta_{\text{min}}}{8L^{2}_{12}},\frac{1}{4\beta_{\max}}\}$, $\underline{f}=\min_{(x,y)\in X\times Y}f(x,y)$, $\beta_{\text{min}}={l^{1}_{11}}+{l^{1}_{12}}+\frac{32(l_{12}^{1})^{2}(l^{1}_{11}+l^{1}_{22})}{\mu_{1}^{2}}$ and $\beta_{\text{max}}=2L_{11}+2{L_{12}}+\frac{2^{9}L_{12}^{2}(L_{12}+L_{22})}{\mu^{2}}$.
@@
Let $d^{(k)}_{1}={\min\left\{\frac{2\gamma_{k}-\frac{l^{k}_{12}}{2}}{2\gamma_{k}^{2}},\frac{\beta_{k}-\frac{l^{k}_{11}}{2}-\frac{l^{k}_{12}}{2}-\frac{16(l^{k}_{12})^{2}\gamma_{k-1}}{\mu_{k-1}\mu_{k}}}{\beta_{k}^{2}+2L_{12}^{2}}\right\}}>0$. Multiplying both sides of (34) by $d^{(k)}_{1}$, and using (22) in Lemma 5, we have
@@
$F_{T(\varepsilon)+1}$ $=f_{T(\varepsilon)+1}+S_{T(\varepsilon)+1}-\frac{2\gamma_{T(\varepsilon)}^{2}}{\mu_{T(\varepsilon)}}\mathcal{D}^{2}_{y}-\frac{7}{2}\gamma_{T(\varepsilon)+1}||y_{T(\varepsilon)+1}-y_{T(\varepsilon)}||^{2}$
$\geq\underline{f}-\left(14L_{12}+\frac{128L_{12}}{\mu}\right)\mathcal{D}^{2}_{y}=\underline{F},$ (38)
@@
In PF-AGP-NSC algorithm, the  total number of backtracking step is upper bounded by  $\sum\limits_{k=2}^{T(\varepsilon)}(\log_{2}\left(\frac{l^{k}_{11}}{l_{11}^{k-1}}\right)+\log_{2}\left(\frac{l^{k}_{12}}{l_{12}^{k-1}}\right)+\log_{2}\left(\frac{l^{k}_{22}}{l_{22}^{k-1}}\right)+\log_{2}\left(\frac{\mu_{k-1}}{\mu_{k}}\right))=\log_{2}\left(\frac{2L_{11}}{l_{11}^{1}}\right)+\log_{2}\left(\frac{2L_{12}}{l_{12}^{1}}\right)+\log_{2}\left(\frac{2L_{22}}{l_{22}^{1}}\right)+\log_{2}\left(\frac{2\mu_{1}}{\mu}\right)=O({\log(\kappa}))$. According to Theorem 6,   $d_{1}=O(\frac{1}{L\kappa^{2}})$, $\underline{F}=O(L\kappa)$ , the total number of gradient calls of PF-AGP-NSC to obtain an $\varepsilon$-stationary point that satisfies $\|\nabla G(x_{k},y_{k})\|\leq\varepsilon$ is upper bounded by  $\mathcal{O}(L^{2}\kappa^{3}\varepsilon^{-2})$ under the nonconvex-strongly concave setting, where $\kappa$ is the condition number.
Remark 8
If additional information is available, the iteration complexity of the algorithm can be further improved. To illustrate this point, if $\mu$ and $\underline{S}$ are known in advance, in Appendix A we propose a restarted variant of the PF-AGP-NSC algorithm, named rPF-AGP-NSC, and show that the gradient complexity of the algorithm can be improved to $\mathcal{O}\left(\log(L)L\kappa^{2}\varepsilon^{-2}\right)$, which is better than SGDA-B.